# The Existence of MacWilliams-Type Identities for the Lee, Homogeneous and Subfield Metric

### v2

@@ 3.1. Lee partition for Additive Weights
Figure 1. Illustration of $\mathrm{Comp}^{\mathsf{L}}_{\pi}(\rho)$.
Different to the original MacWilliams identities for the Hamming metric, we aim at enumerating all codewords of the same Lee decomposition $\pi\in\mathbb{D}^{\mathsf{L}}_{p,r,s,n}$. That is, for each $\pi\in\mathbb{D}^{\mathsf{L}}_{p,r,s,n}$, we are interested in
(51) $\mathcal{D}_{\pi}^{\mathsf{L}}(\mathcal{C})\coloneqq\left|\,{P^{\mathsf{L}}_{\pi}\cap\mathcal{C}}\,\right|.$
Furthermore, if $\mathbb{D}^{\mathsf{L}}_{p,r,s,n}=\left\{{\pi^{(0)},\dots,\pi^{(D-1)}}\right\}$, we denote by
(52) $\mathcal{D}^{\mathsf{L}}(\mathcal{C}):=\left(\mathcal{D}_{\pi^{(0)}}^{\mathsf{L}}(\mathcal{C}),\ldots,\mathcal{D}_{\pi^{(D-1)}}^{\mathsf{L}}(\mathcal{C})\right)$
the Lee decomposition enumerator. If the code $\mathcal{C}$ considered is clear from the context, we will simply write $\mathcal{D}^{\mathsf{L}}{}$ and $\mathcal{D}_{\pi^{(i)}}^{\mathsf{L}}{}$, for $i\in\left\{{0,\ldots,D-1}\right\}$.
Note that the Lee partition is always finer than the partition induced by any additive weight $\operatorname{wt}$. This is because, for any weight $\operatorname{wt}$, the weight of each element must be equal to the weight of its opposite. Consequently, the blocks of the partition of $\mathcal{R}^{n}$ induced by $\operatorname{wt}$ contain the blocks of the Lee partition. Thus, for an $n$-tuple $x\in\mathcal{R}^{n}$ we can compute its weight $\operatorname{wt}(x)$ using its Lee decomposition $\pi^{\mathsf{L}}(x)$, i.e.,
@@
By abuse of notation, we call this the $\operatorname{wt}$-weight of $\pi$ and we write $\operatorname{wt}(\pi)$. Thus, given any code $\mathcal{C}\subseteq\mathcal{R}^{n}$ with weight enumerator $A^{\operatorname{wt}}(\mathcal{C})=(A^{\operatorname{wt}}_{0},\ldots,A^{\operatorname{wt}}_{nM_{\mathrm{wt}}})$ as defined in (38) we have
(54) $A_{i}^{\operatorname{wt}}(\mathcal{C})=\sum_{\begin{subarray}{c}\pi\in\mathbb{D}^{\mathsf{L}}_{p,r,s,n}\\ \operatorname{wt}(\pi)=i\end{subarray}}\mathcal{D}_{\pi}^{\mathsf{L}}(\mathcal{C}).$
It is important to note that the Lee decomposition of a codeword gives us information about both the size of the codeword’s support and composition. Additionally, the Lee partition can be interpreted as the orbits of the action of the group {±$\left\{{\pm 1}\right\}$ on the ring $\mathcal{R}$. It has been shown in [19, Theorem 2.6 and Theorem 2.7] that a partition $\mathcal{P}_{H}$ of finite group $G$ defined by the orbits of a subgroup $H\leq\mathrm{Aut}(G)$ is reflexive and allows for a MacWilliams identity. This immediately implies that the MacWilliams-type identities for the Lee decompositions exist for every finite chain ring $\mathcal{R}$ endowed with any additive weight.
In particular, if $\mathcal{R}=\mathbb{Z}/{p^{s}}\mathbb{Z}$, i.e., $r=1$, for all $i\in\left\{{0,\dots,M}\right\}$, the sets $I_{i}^{\mathsf{L}}$ reduce to $I_{i}^{\mathsf{L}}=\left\{{\pm i}\right\}$, $\operatorname{Tr}=\mathrm{id}$, and [19, Theorem 2.7] provides MacWilliams-type identities for Lee-metric codes.
Corollary 26 (MacWilliams-type Identities in the Lee Metric).
For a positive integer $n$, consider a linear code $\mathcal{C}\subseteq(\mathbb{Z}/{p^{s}}\mathbb{Z})^{n}$ and its dual $\mathcal{C}^{\perp}$. For a given Lee decomposition $\rho\in\mathbb{D}_{p,1,s,n}$, the number of codewords in the dual $\mathcal{C}^{\perp}$ of Lee decomposition $\rho$ is given by
(55) $\mathcal{D}_{\rho}^{\mathsf{L}}(\mathcal{C}^{\perp})=\frac{1}{\left|\,{\mathcal{C}}\,\right|}\sum_{\pi\in\mathbb{D}^{\mathsf{L}}_{p,r,s,n}}K^{\mathsf{L}}_{\pi}(\rho)\mathcal{D}_{\pi}^{\mathsf{L}}(\mathcal{C}),$
where $K^{\mathsf{L}}_{\pi}(\rho)$ denotes the value of the Krawtchouk coefficient of $\rho$ with respect to $\pi$ and, for a given $p^{s}$-th root of unity $\xi$, it is given by
(56) $K^{\mathsf{L}}_{\pi}(\rho)\coloneqq\begin{cases}\sum_{t\in\mathrm{Comp}^{\mathsf{L}}_{\pi}(\rho)}\left(\prod_{i=0}^{M}\binom{\pi_{i}}{t_{i0},\ldots,t_{iM}}\prod_{j=1}^{M-1}\left(\xi^{-ij}+\xi^{ij}\right)^{t_{ij}}\xi^{iM}\right)&\text{if }p=2,\\ \sum_{t\in\mathrm{Comp}^{\mathsf{L}}_{\pi}(\rho)}\left(\prod_{i=0}^{M}\binom{\pi_{i}}{t_{i0},\ldots,t_{iM}}\prod_{j=1}^{M}\left(\xi^{-ij}+\xi^{ij}\right)^{t_{ij}}\right)&\text{otherwise.}\\ \end{cases}$
To illustrate the calculations of the identities, let us now give an example for the Lee-metric MacWilliams-type identities.
Example 27.
Consider the code $\mathcal{C}\subseteq(\mathbb{Z}/{9}\mathbb{Z})^{3}$ with generator matrix $G=\begin{pmatrix}3&2&8\end{pmatrix}$. That is,
[…]
(58) $\pi^{(0)}=(3,0,0,0,0),\ \pi^{(1)}=(1,0,0,2,0),\ \pi^{(2)}=(0,1,1,1,0),\ \pi^{(3)}=(0,0,1,1,1)\text{ and }\pi^{(4)}=(0,1,0,1,1).$
(59) $G^{\perp}=\begin{pmatrix}1&0&3\\ 0&1&2\end{pmatrix}.$
(60) $\prod_{i=0}^{M}\binom{\pi_{i}}{t_{i0},\ldots,t_{iM}}\prod_{j=1}^{M}\left(\xi^{-ij}+\xi^{ij}\right)^{t_{ij}}=\binom{\pi_{0}}{t_{00},\ldots,t_{0M}}\left(1+1\right)^{t_{03}}=\binom{3}{2,1}(1+1)=6.$
Consider now $\pi=\pi^{(1)}=(1,0,0,2,0)$. There are two compositions $t,s\in\mathrm{Comp}^{\mathsf{L}}_{\pi}(\rho)$, namely
(61) $t=(t_{0},\ldots,t_{M})$ $:\ t_{0}=(1,0,0,0,0),t_{3}=(1,0,0,1,0)\text{ and }t_{1}=t_{2}=t_{4}=(0,\ldots,0),$
(62) $s=(s_{0},\ldots,s_{M})$ $:\ s_{0}=(0,0,0,1,0),s_{3}=(2,0,0,0,0)\text{ and }s_{1}=s_{2}=s_{4}=(0,\ldots,0).$
Hence, for $\pi=\pi^{(1)}$ we get
(63) $\sum_{t\in\mathrm{Comp}^{\mathsf{L}}_{\pi}(\rho)}$ $\left(\prod_{i=0}^{4}\binom{\pi_{i}}{t_{i0},\ldots,t_{i4}}\prod_{j=1}^{4}\left(\xi^{-ij}+\xi^{ij}\right)^{t_{ij}}\right)$
(64) $=\left(\binom{\pi_{0}}{t_{00},\ldots,t_{04}}\prod_{j=1}^{4}\left(\xi^{0}+\xi^{0}\right)^{t_{0j}}\binom{\pi_{3}}{t_{30},\ldots,t_{34}}\prod_{j=1}^{M}\left(\xi^{-3j}+\xi^{3j}\right)^{t_{3j}}\right.$
(65) $\qquad+\left.\binom{\pi_{0}}{s_{00},\ldots,s_{04}}\prod_{j=1}^{4}\left(\xi^{0}+\xi^{0}\right)^{s_{0j}}\binom{\pi_{3}}{s_{30},\ldots,s_{34}}\prod_{j=1}^{M}\left(\xi^{-3j}+\xi^{3j}\right)^{s_{3j}}\right)$
(66) $=\left(\binom{1}{1,0,0,0,0}\binom{2}{1,0,0,1,0}(\xi^{0}+\xi^{0})^{1}\binom{1}{0,0,0,1,0}\binom{2}{2,0,0,0,0}(\xi^{0}+\xi^{0})^{1}\right)$
(67) $=6.$
For $\pi=\pi^{(2)},\pi^{(3)},\pi^{(4)}$ the computations are similar and yield in each case
since $\sum_{i\in\left\langle{3}\right\rangle}\xi^{i}=0$.
Putting the expressions given in (60), (63) and (68) in the right-hand side of (55) yields
(69) $\frac{1}{\left|\,{\mathcal{C}}\,\right|}\sum_{\pi\in\mathbb{D}^{\mathsf{L}}_{p,r,s,n}}K^{\mathsf{L}}_{\pi}(\rho)\mathcal{D}_{\rho}^{\mathsf{L}}(\mathcal{C})=\frac{1}{9}\left(6\cdot 1+6\cdot 2+0\cdot 2+0\cdot 2+0\cdot 2\right)=2,$
Compared to other partitions (like, for instance, partitioning a code $\mathcal{C}\subseteq\mathcal{R}^{n}$ by the codewords of the same weight), the partition into codewords of the same Lee decomposition is rather fine. The finer a partition is, the more complex the enumerator and the identities become. For a linear Lee-metric code $\mathcal{C}\subseteq(\mathbb{Z}/{p^{s}}\mathbb{Z})^{n}$, however, the Lee partition $\mathcal{P}^{\mathsf{L}}=P^{\mathsf{L}}_{\pi^{(0)}}\;|\;\ldots\;|\;P^{\mathsf{L}}_{\pi^{(D-1)}}$ is the coarsest partition for which MacWilliams-type identities exist and which allows to derive the Lee weight enumerator of $\mathcal{C}$, since for every $c\in P_{\pi^{(i)}}$ we have that $\operatorname{wt}_{\mathsf{L}}(c)=\operatorname{wt}_{\mathsf{L}}(\pi^{(i)})$. In fact, any coarser partition $\mathcal{Q}\geq\mathcal{P}^{\mathsf{L}}$ of $\mathcal{C}$ for which the Krawtchouk coefficient is well-defined (and hence MacWilliams-type identities exist) does not give back the Lee weight enumerator of $\mathcal{C}$. A simple example of this would be the partition into the Hamming weights. These partitions are not of our interest, as the goal is to find MacWilliams-type identities for a partition of the code that is in relation with the weight enumerator of a code.
@@
Even though the MacWilliams identities for the Lee partition are valid for any additive weight over a finite chain ring $\mathcal{R}$, it is not always the most practical partition to choose. Indeed, for other metrics, there exist coarser partitions of the ambient space that still allow for MacWilliams-type identities and allow us to recover the weight enumerator of the code. In the following sections, we focus on the homogeneous metric and the subfield metric, and we introduce, for both metrics respectively, a coarser (and, hence, more practical) partition for which MacWilliams identities exist.
3.2. The Homogeneous Metric
We now focus on the homogeneous metric over $\mathcal{R}$, as defined in Definition 8, with residue field size $q=p^{r}$ and nilpotency index $s>1$. In fact, when $s=1$, $\mathcal{R}$ is a finite field, and the homogeneous metric reduces to the Hamming metric, for which MacWilliams identities exist. On the other hand, if $s>1$, MacWilliams identities for the homogeneous weight enumerator in the classical sense do not exist [18, 44]. Similar to the Lee metric, $n$-tuples with a given homogeneous weight can be decomposed in different ways. A natural choice is the homogeneous decomposition, enumerating the number of entries of $x\in\mathcal{R}^{n}$ of given homogeneous weight. That is, given an $n$-tuple $x\in\mathcal{R}^{n}$, its homogeneous decomposition is defined to be a triple $(\pi^{\mathsf{Hom}}_{0}(x),\pi^{\mathsf{Hom}}_{1}(x),\pi^{\mathsf{Hom}}_{q/(q-1)})$ where for each $i\in\left\{{0,1,\frac{q}{q-1}}\right\}$ we set
@@ 3.2. The Homogeneous Metric
(70) $\pi^{\mathsf{Hom}}_{i}(x):=\left|\,{\left\{{k=1,\ldots,n\;|\;\operatorname{wt}_{\mathsf{Hom}}(x_{k})=i}\right\}}\,\right|.$
@@
Hence, we can define the homogeneous decomposition of a code $\mathcal{C}\subseteq\mathcal{R}^{n}$, partitioning $\mathcal{C}$ into the codewords of the same homogeneous decomposition. We denote the homogeneous decomposition enumerator of $\mathcal{C}$ by $\mathsf{Homdec}_{\mathcal{C}}$. However, this partition does not lead to well-defined Krawtchouk coefficients [18].
One main reason is that, given an element $a$ of homogeneous weight $1$ and a $p^{s}$-th root of unity $\xi$, the sum
@@
(72) $Z:=\left\{{0}\right\},\;U:=\mathcal{R}^{\times},\;S:=\mathcal{S}\setminus\left\{{0}\right\},\;\text{and}\;R:=\mathcal{R}\setminus\left(Z\cup U\cup S\right)=\langle\gamma\rangle\setminus\langle\gamma^{s-1}\rangle.$
@@
(74) $\mathcal{I}^{\mathsf{Hom}}:=Z\;|\;U\;|\;S\;|\;R.$
The homogeneous weight-unit partition $\mathcal{I}^{\mathsf{Hom}}$ of $\mathcal{R}$ allows to define the homogeneous weight-unit decomposition of a tuple $x\in\mathcal{R}^{n}$.
Definition 28.
Let $\mathcal{R}$ be a finite chain ring. We define the homogeneous weight-unit decomposition of an element $x\in\mathcal{R}^{n}$ as a tuple $\pi^{\mathsf{Hom}}(x):=(\pi^{\mathsf{Hom}}_{Z}(x),\pi^{\mathsf{Hom}}_{U}(x),\pi^{\mathsf{Hom}}_{S}(x),\pi^{\mathsf{Hom}}_{R}(x))$, where
(75) $\pi^{\mathsf{Hom}}_{I}(x)=\left|\,{\left\{{k=1,\ldots,n\;|\;x_{k}\in I}\right\}}\,\right|,\text{ for }I\in\left\{{Z,U,S,R}\right\}.$
We denote the set of all homogeneous weight-unit decompositions in $\mathcal{R}^{n}$ by
(76) $\mathbb{D}^{\mathsf{Hom}}_{p,r,s,n}=\left\{{\pi\in\left\{{0,\ldots,n}\right\}^{4}\Big{|}\sum_{I\in\left\{{Z,U,S,R}\right\}}\pi_{I}=n}\right\}.$
Definition 29.
If $\mathbb{D}^{\mathsf{Hom}}_{p,r,s,n}=\left\{{\pi^{(0)},\dots,\pi^{(D-1)}}\right\}$, for some positive integer $D$ is the set of all homogeneous weight-unit decompositions in $\mathcal{R}^{n}$, we define the homogeneous weight-unit partition of $\mathcal{R}^{n}$ as
(77) $\mathcal{P}^{\mathsf{Hom}}:=P^{\mathsf{Hom}}_{\pi^{(0)}}\;|\;P^{\mathsf{Hom}}_{\pi^{(1)}}\;|\;\ldots\;|\;P^{\mathsf{Hom}}_{\pi^{(D-1)}},$
where, for every $i\in\left\{{0,\ldots,D-1}\right\}$,
(78) $P^{\mathsf{Hom}}_{\pi^{(i)}}\coloneqq\left\{{x\in\mathcal{R}^{n}\;|\;\pi^{\mathsf{Hom}}(x)=\pi^{(i)}}\right\}.$
Finally, given two homogeneous weight-unit decompositions $\pi,\rho\in\mathbb{D}^{\mathsf{Hom}}_{p,r,s,n}$, we denote the set of all compositions of $\rho$ with respect to $\pi$ as
(79) $\mathrm{Comp}^{\mathsf{Hom}}_{\pi}(\rho):=\Big{\{}t=(t_{Z},t_{U},t_{S},t_{R})\ \big{|}$ $\ t_{i}=(t_{iZ},\ldots,t_{iR})\in\left\{{0,\ldots,\pi_{i}}\right\}^{4}\text{ with }$
(80) $\sum_{j\in\left\{{Z,U,S,R}\right\}}t_{ij}=\pi_{i}\text{ and }\sum_{i\in\left\{{Z,U,S,R}\right\}}t_{ij}=\rho_{j}\ \forall i,j\Big{\}}.$
Given an $\mathcal{R}$-linear code $\mathcal{C}$, the number of codewords $c\in\mathcal{C}$ of homogeneous weight-unit decomposition $\pi\in\mathbb{D}^{\mathsf{Hom}}_{p,r,s,n}$ is computed as
(81) $\mathcal{D}_{\pi}^{\mathsf{Hom}}(\mathcal{C}):=\left|\,{P^{\mathsf{Hom}}_{\pi}\cap\mathcal{C}}\,\right|.$
Furthermore, note that every $c\in P^{\mathsf{Hom}}_{\pi}$ has the same homogeneous weight given by $\operatorname{wt}_{\mathsf{Hom}}(c)=\pi_{U}+\pi_{R}+\frac{p}{p-1}\pi_{S}$. Hence, the homogeneous weight enumerator $\mathsf{Homwe}_{\mathcal{C}}$ can immediately be derived from the homogeneous weight-unit enumerator. MacWilliams-type identities exists for the homogeneous weight-unit enumerator.
Theorem 30 (MacWilliams-type Identities for the Homogeneous Weight-Unit Enumerator).
Let $\mathcal{R}$ be a finite chain ring with residue field size $q=p^{r}$ and nilpotency index $s$. Consider a linear code $\mathcal{C}\subseteq\mathcal{R}^{n}$ and let $\mathcal{C}^{\perp}$ denote its dual. For a given homogeneous weight-unit decomposition $\rho\in\mathbb{D}^{\mathsf{Hom}}_{p,r,s,n}$, the number of codewords in the dual $\mathcal{C}^{\perp}$ of homogeneous weight-unit decomposition $\rho$ is given by
(82) $\mathcal{D}_{\rho}^{\mathsf{Hom}}(\mathcal{C}^{\perp})=\frac{1}{\left|\,{\mathcal{C}}\,\right|}\sum_{\pi\in\mathbb{D}^{\mathsf{Hom}}_{p,r,s,n}}K^{\mathsf{Hom}}_{\pi}(\rho)\mathcal{D}_{\pi}^{\mathsf{Hom}}(\mathcal{C}).$
Here $K^{\mathsf{Hom}}_{\pi}(\rho)$ denotes the value of the Krawtchouk coefficient of $\rho$ with respect to $\pi$, and for a given $p^{s}$-th root of unity $\xi$ is given by
(83) $K^{\mathsf{Hom}}_{\pi}(\rho)=\sum_{t\in\mathrm{Comp}^{\mathsf{Hom}}_{\pi}(\rho)}\Bigg{(}$ $\prod_{i\in\left\{{Z,U,S,R}\right\}}\binom{\pi_{i}}{t_{iZ},\ldots,t_{iR}}\Bigg{)}(-1)^{t_{US}}(-q^{s-1})^{t_{SU}}\left(q^{s-1}(q-1)\right)^{t_{ZU}}$
(84) $\left(q^{s-1}-q\right)^{t_{ZR}}(q-1)^{t_{ZS}+t_{SS}+t_{SR}+t_{RS}}(-q)^{t_{RR}}\mathbbm{1}_{\left\{{t_{UU}=t_{UR}=t_{RU}=0}\right\}}.$
@@
We start by proving that, for a given $a\in\mathcal{R}^{n}$ with homogeneous weight-unit decomposition $\pi$, the sum $\sum_{x\in P^{\mathsf{Hom}}_{\rho}}\chi_{a}(x)=\sum_{x\in P^{\mathsf{Hom}}_{\rho}}\xi^{\operatorname{Tr}(\left\langle{a,x}\right\rangle)}$ provides the correct expression for the Krawtchouk coefficients, which are indeed well-defined. Note that the summation is made over all $x\in\mathcal{R}^{n}$ with given homogeneous weight-unit decomposition $\rho$. Furthermore, it is only dependent on the homogeneous weight-unit decomposition, but not on the composition of the tuples itself. Therefore, without loss of generality, we assume that the first $\pi_{Z}$ entries of $a$ have weight 00, the next $\pi_{U}$ entries are units. We then find $\pi_{S}$ entries lying in $S=\langle\gamma^{s-1}\rangle\setminus\{0\}$ and finally $\pi_{R}$ entries lying in $R=\left\langle{\gamma}\right\rangle\setminus\left\langle{\gamma^{s-1}}\right\rangle$.
In the first $\pi_{Z}$ positions, an $n$-tuple $x\in\mathcal{R}^{n}$ of homogeneous-weight unit decomposition $\rho$ can have $t_{Zj}<\rho_{j}$ positions lying in $j\in\left\{{Z,U,S,R}\right\}$, where $\sum_{j\in\left\{{Z,U,S,R}\right\}}t_{Zj}=\pi_{Z}$. Repeating the argument iteratively for each $i\in\left\{{Z,U,S,R}\right\}$, we deduce that the number of $x$ satisfying these conditions is then given by the product of multinomial coefficients
(85) $\prod_{i\in\left\{{Z,U,S,R}\right\}}\binom{\pi_{i}}{t_{iZ},\ldots,t_{iR}}.$
Now, for any position $k\in\left\{{1,\ldots,n}\right\}$, for each $I,J\in\left\{{Z,U,S,R}\right\}$ and $a_{k}\in I$, we compute
(86) $\sum_{x_{k}\in J}\xi^{\operatorname{Tr}(a_{k}x_{k})}.$
It is easy to see that if $J=Z$, for $I\in\left\{{Z,U,S,R}\right\}$ and any $a_{k}\in I$, we get
(87) $\sum_{x_{k}\in Z}\xi^{\operatorname{Tr}(a_{k}x_{k})}=\xi^{0}=1.$
@@
Furthermore, note that if $I=Z$ we have $a_{k}=0$. Thus, for every $J\in\left\{{Z,U,S,R}\right\}$ we obtain
(88) $\sum_{x_{k}\in J}\xi^{\operatorname{Tr}(a_{k}x_{k})}=\sum_{y\in J}\xi^{0}=\left|\,{J}\,\right|,$
where the cardinalities $\left|\,{J}\,\right|$ with $J\in\left\{{Z,U,S,R}\right\}$ are given in (73).

### v3

@@ 3.1. Symmetrized Lee Partition for Additive Weights
For any $k\in\left\{{0,\dots,M}\right\}$, $\alpha_{k}$ denotes any element in $I_{k}^{\mathsf{L}}$ and $\binom{\pi_{i}}{t_{i0},\ldots,t_{iM}}$ denotes the multinomial coefficient.
Proof.
Note that, for $a\in P^{\mathsf{L}}_{\pi}$, the sum $\sum_{x\in P^{\mathsf{L}}_{\rho}}\chi_{a}(x)$ runs over all $x\in\mathcal{R}^{n}$ with given Lee decomposition $\rho$ and is only dependent on the Lee decomposition, but not on the composition of the tuples themselves. Therefore, without loss of generality, we assume that the first $\pi_{0}$ entries of $a$ have weight 0, the next $\pi_{1}$ entries equal to $\pm\alpha_{1}$. We then find $\pi_{2}$ entries equals to $\pm\alpha_{2}$ and so on. In the first $\pi_{0}$ positions, an $n$-tuple $x\in\mathcal{R}^{n}$ of Lee decomposition $\rho$ can have $t_{0j}<\rho_{j}$ positions lying in $j\in\left\{{0\dots,M}\right\}$, where $\sum_{j=0}^{M}t_{0j}=\pi_{0}$. Repeating the argument iteratively for each $i\in\left\{{0,\dots,M}\right\}$, we deduce that the number of $x$ satisfying these conditions is then given by the product of multinomial coefficients
(44) $\prod_{i=0}^{M}\binom{\pi_{i}}{t_{i0},\ldots,t_{iM}}$.
The formula follows from the values of the Krawtchouk coefficients with the multiplicities provided by the composition $t\in\mathrm{Comp}^{\mathsf{hom}}_{\pi}(\rho)$. ∎
In particular, if $\mathcal{R}=\mathbb{Z}/{p^{s}}\mathbb{Z}$, i.e., $r=1$, for all $i\in\left\{{0,\dots,M}\right\}$, the sets $I_{i}^{\mathsf{L}}$ reduce to $I_{i}^{\mathsf{L}}=\left\{{\pm i}\right\}$, $\operatorname{Tr}=\mathrm{id}$, and MacWilliams-type identities for the Lee metric codes hold. In this case, we obtain,
[…]
A concrete example to illustrate the calculations of the identities for Lee-metric codes $\mathcal{C}\subseteq(\mathbb{Z}/{p^{s}}\mathbb{Z})^{n}$ is given in A.1.
Compared to other partitions (like, for instance, partitioning a code $\mathcal{C}\subseteq\mathcal{R}^{n}$ by the codewords of the same weight), the partition into codewords of the same Lee decomposition is rather fine. The finer a partition is, the more complex the enumerator and the identities become.
Lemma 25.
Given a finite chain ring $\mathcal{R}$ and a code $\mathcal{C}\subseteq\mathcal{R}^{n}$. The symmetrized Lee partition is the coarsest symmetrized Fourier-reflexive partition that allows retrieving the Lee weight enumerator of $\mathcal{C}$.
Proof.
[…]
Note that the Lee partition is always finer than the partition induced by any weight $\operatorname{wt}$ on $\mathcal{R}$. This is because, for any weight $\operatorname{wt}$, the weight of each element must be equal to the weight of its opposite. Consequently, the blocks of the partition of $\mathcal{R}^{n}$ induced by $\operatorname{wt}$ contain the blocks of the Lee partition. Thus, for an $n$-tuple $x\in\mathcal{R}^{n}$ we can compute its weight $\operatorname{wt}(x)$ using its Lee decomposition $\pi^{\mathsf{L}}(x)$, i.e.,
@@
Hence, the MacWilliams identities for the symmetrized Lee partition are valid for any additive weight over a finite chain ring $\mathcal{R}$. However, the symmetrized Lee partition is not always the most practical partition to choose. Indeed, for other metrics, there exist coarser partitions of the ambient space that still allow for MacWilliams-type identities and allow us to recover the weight enumerator of the code.
3.2. Symmetrized Homogeneous-Unit Partition
We now focus on the homogeneous metric over $\mathcal{R}$, where we assume that $\mathcal{R}$ has nilpotency index $s>1$. In fact, when $s=1$, $\mathcal{R}$ is a finite field, and the homogeneous metric reduces to the Hamming metric for which MacWilliams identities exist.
Similar to the Lee metric, $n$-tuples with a given homogeneous weight can be decomposed in different ways. A natural choice is the symmetrized homogeneous weight partition induced by the homogeneous weight partition on $\mathcal{R}$, classifying the elements $x\in\mathcal{R}^{n}$ of given homogeneous weight. That is, we partition $\mathcal{R}$ into blocks containing all elements of the same homogeneous weight, i.e.,
@@ 3.2. Symmetrized Homogeneous-Unit Partition
(47) $\mathcal{Q}^{\hom}=Q_{0}\;|\;Q_{1}\;|\;Q_{\lambda}\;\text{ where }\;Q_{i}=\left\{{a\in\mathcal{R}\;|\;\operatorname{wt}_{\mathsf{hom}}(a)=i}\right\}.$
@@
Hence, we can define the symmetrized homogeneous weight partition $\mathcal{Q}^{\hom}_{\mathrm{sym},n}$ of a code $\mathcal{C}\subseteq\mathcal{R}^{n}$, partitioning $\mathcal{C}$ into the codewords of the same homogeneous decomposition. However, this partition does not lead to well-defined Krawtchouk coefficients and hence the MacWilliams identities fail [26]. One main reason is that, given an element $a$ of homogeneous weight $1$ and a $p^{s}$-th root of unity $\xi$, the sum
@@
$Z:=\left\{{0}\right\},\;U:=\mathcal{R}^{\times},\;S:=\mathcal{S}\setminus\left\{{0}\right\},\;$
$\text{and}\;R:=\mathcal{R}\setminus\left(Z\cup U\cup S\right)=\langle\gamma\rangle\setminus\langle\gamma^{s-1}\rangle.$
@@
(50) $\mathcal{P}^{\mathsf{hom}}:=Z\;|\;U\;|\;S\;|\;R.$
Let $\mathcal{P}_{\mathrm{sym},n}^{\mathsf{hom}}=P^{\mathsf{hom}}_{\pi^{(0)}}\;|\;\ldots\;|\;P^{\mathsf{hom}}_{\pi^{(D-1)}}$ denote the induced symmetrized homogeneous weight-unit partition on $\mathcal{R}^{n}$, where $\pi^{(i)}\in\mathbb{D}^{\mathsf{hom}}_{n}$, for every $i\in\left\{{0,\ldots,D-1}\right\}$. Furthermore, note that every $c\in P^{\mathsf{hom}}_{\pi}$ has the same homogeneous weight given by $\operatorname{wt}_{\mathsf{hom}}(c)=\pi_{U}+\pi_{R}+\frac{q}{q-1}\pi_{S}$. Hence, the homogeneous weight enumerator $\mathsf{Homwe}_{\mathcal{C}}$ can immediately be derived from the symmetrized homogeneous weight-unit enumerator.
While some Fourier-reflexive partitions over $\mathcal{R}$ naturally arise from a partition of orbits of a subgroup of its automorphism group (e.g., the symmetrized Lee weight partition), this is not generally true for the homogeneous-unit partition $\mathcal{P}^{\mathsf{hom}}$. Recall that the linear isometries preserving the homogeneous weight are given by $\mathrm{Aut}^{\hom}(\mathcal{R})=U^{n}\rtimes S_{n}$.
Proposition 26.
The homogeneous-unit partition $\mathcal{P}^{\mathsf{hom}}$ can be seen as a partition given by the orbits of a subgroup $H\leq\mathrm{Aut}^{\hom}(\mathcal{R})$ if and only if $s\in\left\{{1,2,3}\right\}$.
@@
In the case where $\mathcal{P}^{\mathsf{hom}}$ coincides with the Hamming weight partition (that is, if $s=1$), the partition is clearly induced by a partition of orbits. Additionally, if $s\in\left\{{2,3}\right\}$, $\mathcal{P}^{\mathsf{hom}}$ is given by a partition of the orbits given by the action of $\mathcal{R}^{\times}$ on $\mathcal{R}$.
However, if $s\geq 4$, the situation changes. Indeed, in this case, the set $R$ is given by
(51) $R=\left(\left\langle{p}\right\rangle\cup\cdots\cup\left\langle{p^{s-2}}\right\rangle\right)\setminus\left\langle{p^{s-1}}\right\rangle.$
If, by contradiction, $\mathcal{P}^{\mathsf{hom}}$ does coincide with a partition of an orbit, there would exist an element $r\in R$ such that $R=\left\{{ru\;|\;u\in U}\right\}$. Assume that $r\in\left\langle{p^{i}}\right\rangle\setminus\left\langle{p^{i+1}}\right\rangle$, for some $i\in\{1,\ldots,s-2\}$. Then we observe
(52) $\left\{{ru\;|\;u\in U}\right\}=\left\langle{p^{i}}\right\rangle\setminus\left\langle{p^{i+1}}\right\rangle\neq R,$
which contradicts the assumption. Hence, $\mathcal{P}^{\mathsf{hom}}$ cannot be represented as the partition of orbits. ∎
Although $\mathcal{P}^{\mathsf{hom}}$ (and hence also $\mathcal{P}_{\mathrm{sym},n}^{\mathsf{hom}}$) is not generally representable as a partition of orbits, it is still a Fourier-reflexive partition.
Proposition 27.
The symmetrized homogeneous weight-unit partition $\mathcal{P}_{\mathrm{sym},n}^{\mathsf{hom}}$ is Fourier-reflexive.
Proof.
By [18, Theorem 3.3] it suffices to show that $\mathcal{P}^{\mathsf{hom}}$, as defined in (50), contains $\left\{{0}\right\}$ as a block and is Fourier-reflexive. By the definition of $\mathcal{P}^{\mathsf{hom}}$ the set $\left\{{0}\right\}$ is a block of $\mathcal{P}^{\mathsf{hom}}$. We now show that $\mathcal{P}^{\mathsf{hom}}$ is Fourier-reflexive which holds if and only if $\mathcal{P}^{\mathsf{hom}}$ and $\widehat{\mathcal{P}^{\mathsf{hom}}}$ consist of the same number of blocks. To show that, we identify $\mathcal{R}$ and $\widehat{\mathcal{R}}$ via the isomorphism $a\longmapsto\chi_{a}$, and, for each $a\in\mathcal{R}$ we compute the sums $\sum_{x\in J}\chi_{a}(x)$ for each $J\in\left\{{Z,U,S,R}\right\}$. In what follows, we will set $a\in I$ for a fixed $I\in\left\{{Z,U,S,R}\right\}$,
It is easy to see that if $J=Z$, for $I\in\left\{{Z,U,S,R}\right\}$ and any $a\in I$, we get
(53) $\sum_{x\in Z}\xi^{\operatorname{Tr}(ax)}=\xi^{0}=1.$
@@
Furthermore, note that if $I=Z$ we have $a=0$. Thus, for every $J\in\left\{{Z,U,S,R}\right\}$ we obtain
(54) $\sum_{x\in J}\xi^{\operatorname{Tr}(ax)}=\sum_{y\in J}\xi^{0}=\left|\,{J}\,\right|,$